OpenStudy (anonymous):

Two numbers are randomly selected on a number line numbered 1 through 9. The same number can be chosen twice. What is the probability that both numbers are greater than 6?

OpenStudy (ranga):

Probability of choosing 7, 8 or 9 out of 1 through 9 is: 3/9 = 1/3. Probability of choosing 7, 8, 9 twice is: 1/3 * 1/3 = 1/9.

OpenStudy (anonymous):

thank you